OBSERVATION OF THE CH STRETCH BAND OF CH_{5}^{+}

\underline{E. T. WHITE}, J. TANG, and T. OKA, Department of Chemistry, The University of Chicago, Chicago, IL 60637.

Ab initio calculations predict that protonated methane, CH_{5}^{+}, has three low energy structures that are practically equal in energy and that barriers separating their 120 equivalent minima are low^{a}. Thus CH_{5}^{+} provides a new prototype of spectroscopic specimen in which the five equivalent protons are attached to the central carbon atom^{b}. We have observed, using our difference frequency spectrometer and the velocity modulation technique, a weak but rich spectrum of a carbocation containing one carbon. We believe they are transitions of CH_{5}^{+} due to the formally forbidden CH stretch band based on their plasma chemical behavior and agreement with the conclusion of predissociation spectroscopy of CH_{7}^{+}^{c}. A preliminary observational result will be reported.

aP. R. Schreiner, S. -J., Kim, H. F. Schaefer III, and P. von Ragué Schleyer, J. Chem. Phys. \underline{99}, 3716 (1993)

bP. R. Bunker, J. Mol. Spectrosc. \underline{176}, 297 (1996)

cD. W. Boo and Y. T. Lee, Chem. Phys. Lett. \underline{211}, 358 (1993)